### v1

@@ I Introduction
It is worth noting that, in industrial setups, anomalies are not generated by Gaussian processes but are the outcome of specific, often predictable, issues during the production process. Consequently, the anomalous samples are not randomly distributed outside the nominal distribution; they can be modeled as a mixture of Gaussian distributions in the feature space instead. While general, unpredictable anomalies can still happen. Expert operators can easily define the main problems they can expect from the manufacturing process, such as which kind of defects, in which locations, and how often they expect to appear. Thus, generative AI can represent a powerful tool for SDD, with defect image generation emerging as a promising approach to enhance detector performance. Specifically, in recent years, Denoising Diffusion Probabilistic Models (DDPMs) [10] received significant attention as a powerful class of generative models.
@@ II Related Work
Thus, many recent approaches adopt a “clean” setup, where the training set consists of only nominal samples. Two strategies can be adopted in clean setups: model fitting and image generation. Model fitting approaches aim at generating an accurate model of the nominal distribution, considering an outlier in every sample with a likelihood lower than –or a distance from the nominal prototype higher than– a predefined threshold [7, 14]. On the contrary, data augmentation approaches leverage generative methods to synthesize images of defects and use these images as positive samples for training a supervised model. Specifically, this work focuses on generation-based data augmentation under clean setups.
@@ III-A Multimodal diffusion-based image generation
DDPMs [19, 10] are a class of deep latent variable models that work by modeling the joint distribution of the data over a Markovian inference process. This process consists of small perturbations of the data with a variance-preserving property [20], such that the limit distribution after the diffusion process is approximately identical to a known prior distribution. Starting with samples from the prior, a reverse diffusion process is learned by gradual denoising the sample to resemble the initial data by the end of the procedure.
@@
In addition, we leveraged the natural ability of DDPMs to incorporate multimodal conditioning in the generation process, taking inspiration from [21, 22, 23, 24]. Specifically, we use as textual descriptions a prompt and a negative prompt, i.e., a prompt which guides the image generation “away” from its concepts of the desired output, resulting in high-quality images that comply with the given descriptions [25, 26, 27].
@@ III-B The DIAG pipeline
To generate an anomalous image $i_{a}$, the process starts by sampling a random negative image, an anomaly description, and a mask, forming the triplet $(i_{n},d_{a},m_{a})$. These pieces of information will then be fed to a text-conditioned LDM to perform inpainting on image $i_{n}$ using the mask $m_{a}$.
@@ IV-C Quantitative results
Here, we emulate the situation where no original positive samples are available in the training set. This scenario makes generating augmented positive samples necessary and restricts the users to augmentation procedures that do not rely on positive images. To do this, we build the set of augmented anomalous images $I_{a}$ by generating $N_{aug}$ augmented positive samples with different pipelines, i.e., MemSeg [15], In&Out [18] and DIAG. Then, we train the ResNet-50 model on a dataset that includes the original negative samples $I_{n}$ and the augmented positive samples $I_{a}$. Finally, we evaluate the model on the original test set.
@@ IV-D Qualitative results
In addition, we provide a numeric evaluation of the similarity between the generated images and the real ones by employing FID [29]. It is worth noting that due to the limited number of anomalous images in the original dataset, we are forced to calculate FID on a different network layer, precisely the second max pooling layer. This is a common procedure in cases where the number of images is low, as the calculation requires the number of samples (images) to be higher than the number of features. Furthermore, it does not change the general behavior of the metric but changes the magnitude of the values obtained, making it an important detail for future comparisons. In the specific case of KSDD2, we choose the first and second max-pooling layers with 64 and 192 features, respectively. Specifically, we compare the images generated with MemSeg, In&Out, and DIAG with the ones available in the KSDD2 dataset and compute the FID scores between the positive images of the KSDD2 and the previously mentioned sets of augmented images.
@@ V Conclusions
This work presents DIAG, a novel data augmentation pipeline that leverages language conditioned Latent Diffusion Models to produce training-free positive images for enhancing the performance of a surface defect detection model. We introduced domain experts in the generation pipeline, asking them to describe with textual prompts how a defect should look and where it can be localized. Then, we adopt a pre-trained LDM to generate defective images and train a binary classifier for isolating the anomalous images. We focus our experiments on the KSDD2 dataset and establish ourselves as the new state-of-the-art data augmentation pipeline, surpassing previous approaches in both the zero-shot and full-shot data augmentation scenarios with an AP of .801 and .924, respectively. These results highlight the potential of in-distribution data augmentation in the anomaly detection field, where training-free generative model pipelines such as DIAG can provide meaningful data for downstream classification, making them appealing solutions in scenarios where real anomalous data is difficult to collect or unavailable. These promising results promote further exploration across various datasets, particularly investigating how robust the image generation is compared to noisy textual prompts.

### v2

@@ I Introduction
It is worth noting that, in industrial setups, anomalies are not generated by Gaussian processes but are the outcome of specific, often predictable, issues during the production process. Consequently, the anomalous samples are not randomly distributed outside the nominal distribution; they can be modeled as a mixture of Gaussian distributions in the feature space instead. Expert operators can easily define the main problems they can expect from the manufacturing process, such as which kind of defects, in which locations, and how often they expect them to appear. Thus, generative AI can represent a powerful tool for SDD, with defect image generation emerging as a promising approach to enhance detector performance. Specifically, in recent years, Denoising Diffusion Probabilistic Models (DDPMs) [10] received significant attention as a powerful class of generative models.
@@ II Related Work
Thus, many recent approaches adopt a “clean” setup, where the training set consists of only nominal samples. Two strategies can be adopted in clean setups: model fitting and image generation. Model fitting approaches aim at generating an accurate model of the nominal distribution, considering an outlier every sample with a likelihood lower than –or a distance from the nominal prototype higher than– a predefined threshold [7, 14]. On the contrary, data augmentation approaches leverage generative methods to synthesize images of defects and use these images as positive samples for training a supervised model. Specifically, this work focuses on generation-based data augmentation under clean setups.
@@ III-A Multimodal diffusion-based image generation
DDPMs [19, 10] are a class of deep latent variable models that work by modeling the joint distribution of the data over a Markovian inference process. This process consists of small perturbations of the data with a variance-preserving property [20], such that the limit distribution after the diffusion process is approximately identical to a known prior distribution. Starting with samples from the prior, a reverse diffusion process is learned by gradually denoising the sample to resemble the initial data by the end of the procedure.
@@
In addition, we leveraged the natural ability of DDPMs to incorporate multimodal conditioning in the generation process, taking inspiration from [21, 22, 23, 24]. Specifically, we will use prompts, i.e., textual descriptions of the anomaly, and negative prompts, i.e., prompts that guide the image generation “away” from its concepts. This results in high-quality images that comply with the given descriptions [25, 26, 27].
@@ III-B The DIAG pipeline
To generate an anomalous image $i_{a}$, the process starts by sampling a random negative image, an anomaly description, and a mask, forming the triplet $(i_{n},d_{a},m_{a})$. Instead of directly operating on the image pixels using DDPM, we use a Latent Diffusion Model (LDM) to work in a lower-dimensional latent space [22]. Thus, the above information will be fed to a text-conditioned LDM to perform inpainting on image $i_{n}$ using the mask $m_{a}$.
@@ IV-C Quantitative results
Here, we emulate the situation where no original positive samples are available in the training set. This scenario makes generating augmented positive samples necessary and restricts the users to augmentation procedures that do not rely on positive images. To do this, we build the set of augmented anomalous images $I_{a}$ by generating $N_{aug}$ augmented positive samples with different pipelines, i.e., MemSeg [15], In&Out [18] and DIAG. Then, we train a ResNet-50 on a dataset that includes the original negative samples $I_{n}$ and the augmented positive samples $I_{a}$. Finally, we evaluate the model on the original test set.
@@ IV-D Qualitative results
In addition, we provide a numeric evaluation of the similarity between the generated images and the real ones by employing FID [29]. It is worth noting that due to the limited number of anomalous images in the original dataset, we are forced to calculate FID on a different network layer, precisely the second max pooling layer. This is a common procedure in cases where the number of images is low, as the calculation requires the number of samples (images) to be higher than the number of features. Note that this only changes the magnitude of the values obtained, not the metric’s general behavior. In the specific case of KSDD2, we choose the first and second max-pooling layers with 64 and 192 features, respectively. Specifically, we compare the images generated with MemSeg, In&Out, and DIAG with the ones available in the KSDD2 dataset and compute the FID scores between the positive images of the KSDD2 and the previously mentioned sets of augmented images.
@@ V Conclusions
This work presents DIAG, a novel data augmentation pipeline that leverages language-conditioned Latent Diffusion Models to produce training-free positive images for enhancing the performance of a surface defect detection model. We introduced domain experts in the generation pipeline, asking them to describe with textual prompts how a defect should look and where it can be localized. Then, we adopt a pre-trained LDM to generate defective images and train a binary classifier for isolating the anomalous images. We focus our experiments on the KSDD2 dataset and establish ourselves as the new state-of-the-art data augmentation pipeline, surpassing previous approaches in both the zero-shot and full-shot data augmentation scenarios with an AP of .801 and .924, respectively. These results highlight the potential of in-distribution data augmentation in the anomaly detection field, where training-free generative model pipelines such as DIAG can provide meaningful data for downstream classification, making them appealing solutions in scenarios where real anomalous data is difficult to collect or unavailable. These promising results promote further exploration across various datasets, particularly investigating how robust the image generation is compared to noisy textual prompts.